OpenStudy (anonymous):

Find the common ratio of the geometric sequence: 1.512, 7.56, 37.8, 189, 945, ....

OpenStudy (anonymous):

the comman ratio is : 5

OpenStudy (anonymous):

:5?

OpenStudy (anonymous):

1:5 or 0:5 ?

OpenStudy (anonymous):

if you divide the first term on second term u will get the comman for example 1.512/7.56 = 5

OpenStudy (anonymous):

r (common ratio) = second term /first term

OpenStudy (anonymous):

r = 7.56/1.512 =5

OpenStudy (anonymous):

sorry second term on first term !!

OpenStudy (anonymous):

37.8 / 7.56 = 945 / 189 = 5